OpenStudy (anonymous):

Please help! Correct answer gets a medal!! There are 9 items at the salad bar. How many ways can Gwen select 3 items from the salad bar? A. 3 B. 27 C. 84 D. 504

OpenStudy (solomonzelman):

I think you basically asked,\[9 ~~~nCr~~~3\]

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

would it be like 9! divided by 3! or?

OpenStudy (solomonzelman):

nCr = n!/(r!(n - r)!)

OpenStudy (solomonzelman):

\[\huge\color{red}{ nCr=\frac{9!}{3!(9-3)!} }\]

OpenStudy (anonymous):

so its C?

OpenStudy (anonymous):

84

OpenStudy (solomonzelman):

\[9-3=6~~~~~~~so,\]\[\color{red}{ \frac{9!}{3! \times 6!} }\] \[9!=6! \times 7 \times 8 \times 9\] and 6! cancels. YOU HAVE LEFT.\[\color{red}{ \frac{7 \times 8 \times 9}{1 \times 2 \times 3} }\]

OpenStudy (solomonzelman):

\[7 \times 4 \times 3=84\]

OpenStudy (solomonzelman):

You are correct, and fast!